OpenStudy (anonymous):

WILL FAN! Equation A: y = 2x + 4 Equation B: y = 3x + 1

OpenStudy (anonymous):

\[\text{Solve}[\{ y=2 x+4, y=3 x+1 \}] \]\[{x = 3, y = 10} \]

OpenStudy (anonymous):

ohhhhhh thank you!!

OpenStudy (anonymous):

{2 x + 4 = 3 x + 1} // Solve x = 3